OpenStudy (anonymous):

A circular arc of length 14 feet subtends a central angle of 25 degrees. Find the radius of the circle in feet

OpenStudy (anonymous):

is the answer approx 32?

OpenStudy (anonymous):

how did you work the problem

OpenStudy (anonymous):

ok, firstly a circle has 360 degrees

OpenStudy (anonymous):

so you need to multiply 14 by 360/25

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

to find a circumference

OpenStudy (anonymous):

we are supposed to leave pi in the answer

OpenStudy (anonymous):

circumference=2*pi*r

OpenStudy (anonymous):



OpenStudy (anonymous):

so the circumference=14*360/25=2*pi*r

OpenStudy (anonymous):

so by this i found r

OpenStudy (anonymous):

so 201.6=2pi*r

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then divide by 2pi? i keep getting 100.8pi but it says its the wrong answer

OpenStudy (anonymous):

ok i got it i had to put32.1 THanks!